OpenStudy (anonymous):

y-intercept of -2 and x-intercept of 5

OpenStudy (anonymous):

y=mx-2 0=5m-2 m=2/5 y=.4x-2

OpenStudy (anonymous):

Wrong!

OpenStudy (anonymous):

Seriosly!

OpenStudy (anonymous):

Who teached math!